# DynamicID: Zero-Shot Multi-ID Image Personalization with Flexible Facial Editability

### v1

@@ Abstract
Recent advancements in text-to-image generation have spurred interest in personalized human image generation, which aims to create novel images featuring specific human identities as reference images indicate. Although existing methods achieve high-fidelity identity preservation, they often struggle with limited multi-ID usability and inadequate facial editability. We present DynamicID, a tuning-free framework supported by a dual-stage training paradigm that inherently facilitates both single-ID and multi-ID personalized generation with high fidelity and flexible facial editability. Our key innovations include: 1) Semantic-Activated Attention (SAA), which employs query-level activation gating to minimize disruption to the original model when injecting ID features and achieve multi-ID personalization without requiring multi-ID samples during training. 2) Identity-Motion Reconfigurator (IMR), which leverages contrastive learning to effectively disentangle and re-entangle facial motion and identity features, thereby enabling flexible facial editing. Additionally, we have developed a curated VariFace-10k facial dataset, comprising 10k unique individuals, each represented by 35 distinct facial images. Experimental results demonstrate that DynamicID outperforms state-of-the-art methods in identity fidelity, facial editability, and multi-ID personalization capability.
@@
footnotetext: * Corresponding Author
@@ 1 Introduction
Previous research in personalized human image generation has achieved substantial progress through various approaches [1, 51, 8, 4, 6, 14, 37, 73]. Early methodologies [56, 16, 57, 34] demonstrated exceptional high-fidelity outcomes but required subject-specific retraining, imposing significant computational burdens and temporal constraints. To address these limitations, subsequent tuning-free paradigms [41, 70, 81, 72, 28, 78] employed pre-trained face encoders coupled with feature injection modules. These frameworks extract facial features through the encoder and subsequently integrate them into the generation process via the injection module. While these parameter-efficient methods eliminate retraining requirements and improve practical deployment capabilities, two critical limitations persist, as shown in Fig. 2: Constraints in Multi-ID Generation: In scenarios involving multiple identities, numerous approaches struggle with the challenge of identity blending [77]. While various mitigation strategies have emerged, these methods are built upon frameworks tailored for the single-ID contexts, which inherently leads to a compromise in model performance. Insufficient Facial Attribute Editability: Current methodologies lack explicit training models to disentangle identity features (e.g., facial structure and skin texture) from motion features (e.g., expression and orientation). This intrinsic entanglement within facial latent representations fundamentally restricts the flexible manipulation of facial attributes while preserving identity fidelity.
To overcome these limitations, we propose DynamicID, a tuning-free method that inherently supports both single-ID and multi-ID personalized image generation, ensuring high identity fidelity and flexible facial editability. DynamicID comprises two novel designs: 1) Semantic-Activated Attention (SAA) and 2) Identity-Motion Reconfigurator (IMR). The SAA, as an innovative feature injection mechanism, dynamically modulates the activation levels of distinct image latent queries based on their semantic relevance to the facial feature keys.
Figure 2: Visual comparison on Multi-ID Personalization. Current methods demonstrate constraints in multi-ID generation and unsatisfying editability, while our method enables flexible facial editability as well as high image quality in such scenarios.
@@
This mechanism ensures that the original model’s behavior remains undisturbed during the injection of ID features and allows for free manipulation of activation levels to enable zero-shot layout control and multi-ID personalized image generation. The IMR, as a trainable feature transformer, learns to effectively disentangle and re-entangle identity and motion features of facial latent representations. It serves as an intermediary by capturing a face encoder’s output, modifying features, and delivering the modified output to the SAA, thereby enabling high fidelity and flexible facial editability.
@@
In the realm of the training paradigm, end-to-end training serves as a simple and straightforward approach. Yet the demand for large-scale specialized datasets with multi-ID concurrence and free-form facial attributes makes it practically unfeasible. To overcome this challenge, we introduce a task-decoupling training paradigm, characterized by a strategic bifurcation into two distinct stages aimed at separately training the SAA and the IMR. In the initial stage, we concurrently train the SAA and a face encoder, leveraging the innovative mechanism of the SAA, which allows this stage to be conducted using images containing only one single individual. Subsequently, our focus shifts to training the IMR which requires only a facial dataset where each individual is represented by multiple distinct images. To further augment the training of the IMR, we have meticulously curated the VariFace-10k dataset, which encompasses 10k unique individuals, each represented by 35 distinct facial images, thereby providing a flexible and diverse foundation for model enhancement.
@@
  • We conduct extensive and comprehensive experiments, which confirm DynamicID‘s efficacy and state-of-the-art performance.
@@ 2.1 Text-to-Image Models
The field of text-to-image generation has experienced significant advancements, driven by innovative techniques such as Generative Adversarial Networks (GANs) [30, 76], auto-regressive models [53, 65], and diffusion models [25, 54]. Among these techniques, diffusion models have emerged as the predominant approach, leveraging diffusion processes to iteratively transform noise into coherent images that align with textual descriptions [50, 47, 48, 64, 15]. In line with current trends, our methodology is built upon the Stable Diffusion model.
@@ 2.2 Personalized Human Image Generation
Tuning-free single-ID personalization: A vast amount of tuning-free studies have concentrated on the single-ID contexts [5, 10, 9, 29, 43], which typically utilize a face encoder to extract facial features from reference images and an injection module to integrate these features into the generative process. Current approaches can be systematically categorized into three distinct types based on the injection mechanism: text-based methods [66, 62, 7, 71], network-based methods [21, 81], and attention-based methods [42, 75, 17, 20, 74]. To illustrate, PhotoMaker [40] extracted facial features and merged them with the corresponding text token within the text embedding space, achieving semantic consistency though compromising identity fidelity. Conversely, InstantID [70] adapts ControlNet [80] architecture for high-fidelity generation but lacks facial editing granularity. Improving upon this, CapHuman [41] enhances facial editability via 3D Morphable Face Model [38] integration at the cost of increased computational complexity. As the pioneering attention-based method, IP-Adapter [79] introduces a novel decoupled cross-attention mechanism for facial identity integration, achieving superior fidelity while sharing InstantID’s limitation in granular facial control. Inspired by this, the $W_{+}$ adapter [39] leveraged the StyleGAN [32] encoder to extract features for expression control. However, using this encoder reduced fidelity and limited control over facial orientation. Moreover, these attention-based methods disrupted the model’s original behavior during the generative process due to the coarse-grained nature of cross-attention integration.
@@
Tuning-free multi-ID personalization: Multi-ID contexts remain significantly understudied compared to single-ID contexts, presenting a unique technical challenge: identity blending [77], where facial features from different reference subjects intermingle during generation. Recent studies [49, 69, 13, 19] have proposed several mitigation strategies. Some works build on text-based methods originally tailored for single-ID contexts. For example, FastComposer [77] addressed identity blending through localized cross-attention maps, while Face-diffuser [72] combined dual specialized diffusion models with Saliency-adaptive Noise Fusion, simultaneously preventing the intermingling of features and enhancing image quality. Other works build on attention-based methods. UniPortrait [24] introduced an ID routing module, enabling precise spatial allocation of facial features during synthesis. InstantFamily [33] and MagicID [13] incorporated a masked cross-attention mechanism to prevent feature leakage, coupled with a ControlNet integration to maintain positional correspondence between generated identities and their masked attention maps. Although prior works have addressed identity blending challenges, their reliance on existing frameworks tailored for single-ID context has inadvertently compromised core model functionalities while failing to achieve precise facial editing.
@@ 3 Methodology
We present a novel unified framework that enables both single-ID and multi-ID personalized generation while achieving high identity fidelity and flexible facial editability. Our framework integrates three components: (1) a face encoder extracting facial features from reference images, (2) the Identity-Motion Reconfigurator (IMR) operating in latent space for feature-level editing, and (3) the Semantic-Activated Attention (SAA) effectively injecting processed features into text-to-image models. To address the challenge of requiring large-scale specialized datasets, we introduce the task-decoupled training paradigm structured as an anchoring stage and a reconfiguration stage. As visualized in Fig. 3, the initial anchoring stage establishes fundamental personalized capabilities by simultaneously optimizing the face encoder and the SAA. The subsequent reconfiguration stage specializes in training the IMR for flexible facial editability. Section 3.1 provides a comprehensive analysis of the SAA alongside implementation details of the anchoring stage. Section 3.2 systematically elaborates on the IMR architecture and the reconfiguration stage methodology. Furthermore, we contribute the VariFace-10k dataset in Section 3.3 to support IMR training through diversified facial variations.
@@ 3.1 Semantic-Activated Attention (SAA)
The predominant technique for feature injection in personalized generation is the cross-attention mechanism. However, the inherent softmax function introduces an implicit inductive bias that forces each query to distribute a fixed total attention mass across all keys, even in the absence of semantic relevance between the query and any of the keys, which potentially causes perturbations to the model’s original behavior and identity blending [77] in multi-ID personalization. To overcome these limitations, we propose the Semantic-Activated Attention (SAA). As shown in Fig. 4, given the latent image representation $z$ and facial features $c_{f}$ extracted from the reference image, SAA is formally defined through the following operations:
@@
where $Q=zW_{q}$, $K=c_{f}W_{k}$, and $V=c_{f}W_{v}$ represent the query, key, and value matrices of the attention mechanism, respectively. The matrix $J\in\mathbb{R}^{k\times 1}$ denotes an all-ones column vector $\begin{bmatrix}1&1&\cdots&1\end{bmatrix}^{\top}$ with length matching the number of keys, $k$. The $\mathrm{Expand}(\cdot)$ operator broadcasts the weight vector $w$ along the feature dimension to match the dimension of the attention output matrix. The $\mathrm{Norm}(\cdot)$ operator applies min-max normalization to scale its input values to the [0, 1] range. The activation weights, derived from the interactions between queries and keys, effectively indicate the extent to which each query should be associated with the reference facial information, as verified in Fig. 7. Intuitively, SAA enhances ID features integration by ensuring that queries corresponding to the facial region receive strong activation (with their corresponding activation weight approaching 1) to incorporate reference facial information, while queries corresponding to the background are suppressed (with their corresponding activation weight approaching 0) according to their inherent irrelevance to the reference face. Additionally, queries corresponding to the body are moderately activated to maintain the internal coherence of the image.
@@
Context Decoupling: The novel SAA suppresses queries originating from non-facial regions during inference, ensuring that the generation process in these areas remains unaffected by extraneous facial information as shown in Fig. 8(a). This decoupling of context allows the model to maintain remarkable consistency in background and character behaviors across different reference identities.
Layout Control: The novel SAA enables parametric modulation of activation weights during early denoising steps to achieve layout control as shown in Fig. 8(b). This controlled intervention, scaled by parameter $\beta$, induces the semantic transition of targeted regions toward facial features. Global coherence is maintained through multi-level query activation in subsequent denoising stages, which demonstrates superiority over binary activation approaches. The intervention is formally expressed as:
@@
where $M$ represents the user-specified mask region, with the hyperparameters $\alpha$ and $\beta$ governing the temporal scope and intervention intensity, respectively.
@@
Notably, besides its versatility, the SAA framework maintains remarkable training efficiency, requiring only the single-ID dataset which is readily accessible. For the face encoder, we employ two pre-trained feature extractors to independently capture global and local facial features, followed by a projection module that synthesizes these features into a unified facial latent representation. In the anchoring stage, we freeze the parameters of the pre-trained extractors, training only the projection module and SAA, using the same training objective as the original diffusion process:
@@ 3.2 Identity-Motion Reconfigurator (IMR)
To achieve personalized generation with flexible and fine-grained facial editability, we introduce the Identity-Motion Reconfigurator (IMR) within the interaction space of the face encoder and the SAA. The IMR comprises two components: DisentangleNet $\boldsymbol{\phi_{1}}$ and EntangleNet $\boldsymbol{\phi_{2}}$, which strategically utilize face prompts to extract facial motion characteristics and incorporate subtle perturbed facial landmarks to enhance the accuracy of orientation angle estimation. The face prompts are encoded into $f_{p}\in\mathbb{R}^{1\times d_{p}}$ through a text encoder, while landmarks are transformed into $f_{l}\in\mathbb{R}^{1\times d_{l}}$ via a keypoint encoder. These representations are subsequently projected into $c$ tokens using a Multi-layer Perceptron (MLP), expressed as:
@@
DisentangleNet, leveraging the facial features $\xi_{\textit{src}}$ extracted by the face encoder trained during the anchoring stage, alongside the corresponding motion features $\psi_{\textit{src}}$, predicts the posterior distribution of the identity features $p(\zeta|\xi_{\textit{src}},\psi_{\textit{src}})$. Subsequently, EntangleNet forecasts the distribution of the target facial features $p(\xi_{\textit{pred}})$ by integrating the identity features $\zeta$ with the target motion features $\psi_{\textit{tgt}}$.
@@
Consequently, by disentangling and re-entangling facial features, the IMR captures the feature output of the face encoder, effectively modifies its motion features towards the desired direction, and then provides these modified features to the SAA. Thereby, DynamicID facilitates fine-grained and flexible control of the facial motion while preserving identity fidelity.

### v2

@@ Abstract
Recent advances in text-to-image generation have driven interest in generating personalized human images that depict specific identities from reference images. Although existing methods achieve high-fidelity identity preservation, they are generally limited to single-ID scenarios and offer insufficient facial editability. We present DynamicID, a tuning-free framework that inherently facilitates both single-ID and multi-ID personalized generation with high fidelity and flexible facial editability. Our key innovations include: 1) Semantic-Activated Attention (SAA), which employs query-level activation gating to minimize disruption to the base model when injecting ID features and achieve multi-ID personalization without requiring multi-ID samples during training. 2) Identity-Motion Reconfigurator (IMR), which applies feature-space manipulation to effectively disentangle and reconfigure facial motion and identity features, supporting flexible facial editing. 3) a task-decoupled training paradigm that reduces data dependency, together with VariFace-10k, a curated dataset of 10,000 unique individuals, each represented by 35 distinct facial images. Experimental results demonstrate that DynamicID outperforms state-of-the-art methods in identity fidelity, facial editability, and multi-ID personalization capability. Our code is available at https://github.com/ByteCat-bot/DynamicID.
@@
footnotetext: Equal contributors.  Corresponding Author
@@ 1 Introduction
Previous research in personalized human image generation has achieved substantial progress through various approaches [1, 52, 8, 4, 6, 14, 38, 74]. Early methodologies [57, 16, 58, 35] demonstrated exceptional high-fidelity outcomes but required subject-specific retraining, imposing significant computational burdens and temporal constraints. To address these limitations, subsequent tuning-free paradigms [42, 71, 82, 73, 28, 79] employed pre-trained face encoders coupled with feature injection modules. These frameworks extract facial features through the encoder and subsequently integrate them into the generation process via the injection module. While these parameter-efficient methods eliminate retraining requirements and improve practical deployment capabilities, two critical limitations persist: Constraints in Multi-ID Generation: In scenarios involving multiple identities, numerous approaches struggle with the challenge of identity blending [78]. While various mitigation strategies have emerged, these methods are built upon frameworks tailored for the single-ID contexts, which inherently compromises model performance. Insufficient Facial Attribute Editability: Current methodologies lack explicit training models to disentangle identity features (e.g., facial structure and skin texture) from motion features (e.g., expression and orientation). This intrinsic entanglement within facial latent representations fundamentally restricts the flexible manipulation of facial attributes while preserving identity fidelity.
@@
To overcome these limitations, we propose DynamicID, a tuning-free method that inherently supports single-ID and multi-ID personalized image generation, ensuring high identity fidelity and flexible facial editability. DynamicID comprises two novel designs: 1) Semantic-Activated Attention (SAA) and 2) Identity-Motion Reconfigurator (IMR). As an innovative feature injection mechanism, the SAA dynamically modulates the activation levels of distinct image latent queries based on their semantic relevance to the facial feature keys. This mechanism ensures that the original model’s behavior remains undisturbed during the injection of ID features and allows for free manipulation of activation levels to enable zero-shot layout control and multi-ID personalized image generation. The IMR, as a trainable feature transformer, learns to effectively disentangle and reconfigure identity and motion features of facial latent representations. It serves as an intermediary by capturing a face encoder’s output, modifying features, and delivering the modified output to the SAA, thereby enabling high fidelity and flexible facial editability.
@@
In the realm of the training paradigm, end-to-end training serves as a straightforward approach. Yet the demand for large-scale specialized datasets with multi-ID concurrence and free-form facial attributes makes it practically unfeasible. To overcome this challenge, we introduce a task-decoupling training paradigm, characterized by a strategic bifurcation into two distinct stages aimed at separately training the SAA and the IMR. In the initial stage, we concurrently train the SAA and a face encoder, leveraging the innovative mechanism of the SAA, which allows this stage to be conducted using images containing only one individual. Subsequently, our focus shifts to training the IMR which requires only a facial dataset where multiple distinct images represent each individual. To further augment the training of the IMR, we have meticulously curated the VariFace-10k dataset, which encompasses 10k unique individuals, each represented by 35 distinct facial images, thereby providing a flexible and diverse foundation for model enhancement.
@@
  • We conduct extensive and comprehensive experiments confirming DynamicID’s efficacy and state-of-the-art performance.
@@ 2.1 Text-to-Image Models
The field of text-to-image generation has experienced significant advancements, driven by innovative techniques such as Generative Adversarial Networks (GANs) [30, 77], auto-regressive models [54, 66], and diffusion models [25, 55]. Among these techniques, diffusion models have emerged as the predominant approach, leveraging diffusion processes to iteratively transform noise into coherent images that align with textual descriptions [51, 48, 49, 65, 15]. Our methodology is built upon the stable diffusion model, which is in line with current trends.
@@ 2.2 Personalized Human Image Generation
Tuning-free single-ID personalization: A vast amount of tuning-free studies have concentrated on the single-ID contexts [5, 10, 9, 29, 44, 68], which typically utilize a face encoder to extract facial features from reference images and an injection module to integrate these features into the generative process. Current approaches can be systematically categorized into three distinct types based on the injection mechanism: text-based methods [67, 63, 7, 72], network-based methods [21, 82], and attention-based methods [43, 76, 17, 20, 75]. To illustrate, PhotoMaker [41] extracted facial features and merged them with the corresponding text token within the text embedding space, achieving semantic consistency though compromising identity fidelity. Conversely, InstantID [71] adapts ControlNet [81] architecture for high-fidelity generation but lacks facial editing granularity. Improving upon this, CapHuman [42] enhances facial editability via 3D Morphable Face Model [39] integration at the cost of increased computational complexity. As the pioneering attention-based method, IP-Adapter [80] introduces a novel decoupled cross-attention mechanism for facial identity integration, achieving superior fidelity while sharing InstantID’s limitation in granular facial control. Inspired by this, the $W_{+}$ adapter [40] leveraged the StyleGAN [32] encoder to extract features for expression control. However, using this encoder reduced fidelity and limited control over facial orientation. Moreover, these attention-based methods disrupted the model’s original behavior during the generative process due to the coarse-grained nature of cross-attention integration.
@@
Tuning-free multi-ID personalization: Multi-ID contexts remain significantly understudied compared to single-ID contexts, presenting a unique technical challenge: identity blending [78], where facial features from different reference subjects intermingle during generation. Recent studies [50, 70, 13, 19, 34] have proposed several mitigation strategies. Some works build on text-based methods originally tailored for single-ID contexts. For example, FastComposer [78] addressed identity blending through localized cross-attention maps, while Face-diffuser [73] combined dual specialized diffusion models with Saliency-adaptive Noise Fusion, simultaneously preventing the intermingling of features and enhancing image quality. Other works build on attention-based methods. UniPortrait [24] introduced an ID routing module, enabling precise spatial allocation of facial features during synthesis. To prevent feature leakage, InstantFamily [33] and MagicID [13] adopted a masked cross-attention mechanism, while StoryMaker [83] introduced a cross-attention loss. All three methods incorporated ControlNet to maintain positional correspondence between generated identities and their masked attention maps. Although prior works have addressed identity blending challenges, their reliance on existing frameworks tailored for single-ID context has inadvertently compromised core model functionalities while failing to achieve precise facial editing.
@@ 3 Methodology
We present a novel unified framework that enables single-ID and multi-ID personalized generation while achieving high identity fidelity and flexible facial editability. Our framework integrates three components: (1) a face encoder extracting facial features from reference images, (2) the Identity-Motion Reconfigurator (IMR) operating in latent space for feature-level editing, and (3) the Semantic-Activated Attention (SAA) effectively injecting processed features into text-to-image models. We introduce the task-decoupled training paradigm structured as anchoring and reconfiguration stages to address the challenge of requiring large-scale specialized datasets. As visualized in Fig. 2, the initial anchoring stage establishes fundamental personalized capabilities by simultaneously optimizing the face encoder and the SAA. The subsequent reconfiguration stage specializes in training the IMR for flexible facial editing. Section 3.1 provides a comprehensive analysis of the SAA alongside implementation details of the anchoring stage. Section 3.2 systematically elaborates on the IMR architecture and the reconfiguration stage methodology. Furthermore, we contribute the VariFace-10k dataset in Section 3.3 to support IMR training through diversified facial variations.
@@ 3.1 Semantic-Activated Attention (SAA)
The predominant technique for feature injection in personalized generation is the cross-attention mechanism. However, the inherent softmax function introduces an implicit inductive bias that forces each query to distribute a fixed total attention mass across keys, even in the absence of semantic relevance between the query and any of the keys. This behavior can lead to two significant issues: 1) perturbations to the model’s original behavior, and 2) identity blending [78] in multi-ID personalization. To overcome these limitations, we propose Semantic-Activated Attention (SAA). As shown in Fig. 3, given the latent image representation $z$ and facial features $c_{f}$ extracted from the reference image, SAA is formally defined through the following operations:
@@
where $Q=zW_{q}$, $K=c_{f}W_{k}$, and $V=c_{f}W_{v}$ represent the attention mechanism’s query, key, and value matrices, respectively. The matrix $J\in\mathbb{R}^{k\times 1}$ denotes an all-ones column vector $\begin{bmatrix}1&1&\cdots&1\end{bmatrix}^{\top}$ with length matching the number of keys, $k$. The $\mathrm{Expand}(\cdot)$ operator broadcasts the weight vector $w$ along the feature dimension to match the dimension of the attention output matrix. The $\mathrm{Norm}(\cdot)$ operator applies min-max normalization to scale its input values to the [0, 1] range. The activation weights, derived from the interactions between queries and keys, effectively indicate the extent to which each query should be associated with the reference facial information, as verified in Fig. 7. Intuitively, SAA enhances ID features integration by ensuring that queries corresponding to the facial region receive strong activation (with their corresponding activation weight approaching 1) to incorporate reference facial information, while queries corresponding to the background are suppressed (with their corresponding activation weight approaching 0) according to their inherent irrelevance to the reference face. Additionally, queries corresponding to the body are moderately activated to maintain the internal coherence of the image.
@@
Context Decoupling: The novel SAA suppresses queries originating from non-facial regions during inference, ensuring that the generation process in these areas remains unaffected by extraneous facial information. As a result, the spatial content and motion dynamics of non-facial regions remain consistent when the reference identity changes, as shown in Fig. 8(a).
Layout Control: The novel SAA enables parametric modulation of activation weights during early denoising steps to achieve layout control as shown in Fig. 8(b). This controlled intervention, scaled by parameter $\beta$, induces the semantic transition of targeted regions toward facial features. Global coherence is maintained through multi-level query activation in subsequent denoising stages, demonstrating superiority over binary activation approaches. The intervention is formally expressed as:
@@
where $M$ represents the user-specified mask region, with the hyperparameters $\alpha$ and $\beta>1$ governing the temporal scope and intervention intensity, respectively.
@@
Notably, besides its versatility, the SAA framework maintains remarkable training efficiency, requiring only the single-ID dataset which is readily accessible. For the face encoder, we employ two pre-trained feature extractors to independently capture global and local facial features, followed by a projection module synthesizing these features into a unified facial latent representation. In the anchoring stage, we freeze the parameters of the pre-trained extractors, training only the projection module and SAA, using the same training objective as the original diffusion process:
@@ 3.2 Identity-Motion Reconfigurator (IMR)
To achieve personalized generation with flexible and fine-grained facial editability, we introduce the Identity-Motion Reconfigurator (IMR) within the interaction space of the face encoder and the SAA. The IMR comprises two components: DisentangleNet $\boldsymbol{\phi_{1}}$ and EntangleNet $\boldsymbol{\phi_{2}}$, which strategically utilize face prompts to extract facial motion characteristics and incorporate subtle perturbed facial landmarks to improve the accuracy of the estimation of the orientation angle. Face prompts are encoded into $f_{p}\in\mathbb{R}^{1\times d_{p}}$ through a text encoder, while landmarks are transformed into $f_{l}\in\mathbb{R}^{1\times d_{l}}$ via a keypoint encoder. These representations are subsequently projected into $c$ tokens using a Multi-layer Perceptron (MLP), expressed as:
@@
DisentangleNet, using the facial features $\xi_{\textit{src}}$ extracted by the face encoder trained during the anchoring stage, together with the corresponding motion features $\psi_{\textit{src}}$, predicts the posterior distribution of the identity features $p(\zeta|\xi_{\textit{src}},\psi_{\textit{src}})$. Subsequently, EntangleNet forecasts the distribution of the target facial features $p(\xi_{\textit{pred}})$ by integrating the identity features $\zeta$ with the target motion features $\psi_{\textit{tgt}}$.
@@
Consequently, by disentangling and reconfiguring facial features, the IMR captures the feature output of the face encoder, effectively modifies its motion features towards the desired direction, and then provides these modified features to the SAA. Thereby, DynamicID facilitates fine-grained and flexible control of facial motion while preserving identity fidelity.